OpenStudy (anonymous):

A conic section has the equation x^2+ y^2+ 12x + 8y = 48. Determine the following: type of conic, domain and range, axes of symmetry, and center. Show your work.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

it is circle .. circle has the form \[x ^{2}+y ^{2}+ax+by+c=0\]

OpenStudy (anonymous):

center (-6,-4) and radius 2

OpenStudy (anonymous):

how did you get 2 for the radius?

OpenStudy (31356):

@crazysingh is correct! :D

OpenStudy (anonymous):

if the circle equation is of the form \[x^{2} + y ^{2} + 2gx + 2fy + c = 0 \] then radius is equal to \[\sqrt{g ^{2}+f ^{2}-c}\]

ganeshie8 (ganeshie8):

radius should be 10 rihgt ?b

OpenStudy (anonymous):

domain is \[x \in (-8,-4)\]

ganeshie8 (ganeshie8):

just check the work again..

OpenStudy (anonymous):

yes radius should be 10

ganeshie8 (ganeshie8):

clearly its a circle,

ganeshie8 (ganeshie8):

domain wud be \(center \pm radius\)

ganeshie8 (ganeshie8):

so domain = \(\large [ -16, 4]\)

OpenStudy (anonymous):

@ganeshie8 : isnt \[domain = x coordinate of center \pm radius\]

ganeshie8 (ganeshie8):

good catch, yes :)

ganeshie8 (ganeshie8):

lets find range

OpenStudy (anonymous):

\[range = y coordinate of center \pm radius\]

OpenStudy (anonymous):

so the domain is not -16, 4?

OpenStudy (anonymous):

domain is represented by \[{ x : -16 \le x \le4 }\]

OpenStudy (anonymous):

ok..and how would we find the range?

OpenStudy (anonymous):

range is represented by \[{y: -14 \le y \le 6 }\]

OpenStudy (anonymous):

ok..what about the axis of symmetry?

OpenStudy (anonymous):

there are infinite numbers of axis of symmetry which passes through center of this circle.. i.e. every line which passes through center of the circle represents axis of symmetry.

OpenStudy (anonymous):

it will have the form : \[(y+6)=m(x+4)\]

OpenStudy (anonymous):

wait..when you say the form, is that how i would put it down as my answer? or do i need to do something with it?

OpenStudy (anonymous):

i guess there is no need to edit..

OpenStudy (anonymous):

ok..thankyou:)

OpenStudy (anonymous):

you should understand question rather than just copying

OpenStudy (anonymous):

i do, and im not "just copying" i was looking at your steps and trying to learn how you did it....

OpenStudy (anonymous):

ok sowiee..